OpenStudy (anonymous):

if 4/3 b=210 solve for b. please help!!!!!

mathslover (mathslover):

is this your question? : \[\large{\frac{4b}{3}=210}\]

OpenStudy (anonymous):

Multiply each side by 3/4.

OpenStudy (chihiroasleaf):

multiply both sides by \[\frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

Multiply by: \[\frac{3}{4}\] Both the sides..

OpenStudy (anonymous):

Why have you closed the question ??

mathslover (mathslover):

\[\large{\color{blue}{\frac{4b}{3}=210}}\] \[\large{\color{red}{\frac{4b}{3}\times 3 = 210 \times 3}}\] \[\large{\color{brown}{\frac{4b}{\cancel{3}^1}\times \cancel{3}^1=210\times 3}}\] \[\large{4b=210\times 3}\] \[\large{\color{green}{\frac{4b}{4}=\frac{210\times 3}{4}}}\] \[\large{\color{orange}{\frac{\cancel{4}^1b}{\cancel{4}^1}=\frac{630}{4}}}\]